# Finite Element Convergence Analysis For Wave Equations With Time-Dependent Coefficients

### v1

@@ 1 Introduction
Wave phenomena are central to a wide range of scientific and engineering applications in acoustics, electromagnetics or elasticity. Traditionally, numerical methods, be they finite element or finite difference methods, have primarily addressed wave propagation in media with material properties varying in space but not in time. Recent advances in material sciences, however, in particular the development of time-modulated metamaterials, have brought renewed attention to wave propagation in passive (or static) media whose properties vary in both space and time. Indeed, these engineered materials enable fine-grained, dynamic control over parameters such as refractive index and impedance, unlocking novel wave-matter interactions with no analogue in static media [1]. From a mathematical perspective, these wave phenomena are modeled by second-order hyperbolic partial differential equations with coefficients that depend explicitly on both space and time, which introduces new analytical and computational challenges.
To this day, the Finite Element Method (FEM) remains probably the most versatile numerical method for the simulation of wave phenomena, as it naturally allows for high-order polynomial approximations, a key feature to efficiently represent the oscillatory nature of waves, but also accomodates complex geometry and discontinuous material interfaces. There is a long history of FEMs for second-order wave equations, starting in the 70’s with the seminal work of Dupont [9] and Baker [3, 2] who first proved optimal $H^{1}$- and $L^{2}$ convergence rates for the classical conforming FEM. Later, convergence was proved in situations of low regularity [12], and more recently in polygons with corner singularities [15]. For non-conforming discontinuous Galerkin (DG) FEMs, convergence for the second-order wave equation was established during the past twenty years (e.g., [16, 11, 10, 5, 8]).
@@ 2 Wave equation in a time-varying medium
where $\mathcal{A}(t)$ is a self-adjoint, coercive and linear second-order elliptic partial differential and $\mathcal{B}(t)$ an arbitrary linear bounded operator. The well-posedness of the variational formulation of (2.5) follows from ([14], Theorem 1.1) under hypotheses about the temporal regularity of $\mathcal{A}(t)$ and $\mathcal{B}(t)$. Although $1/\kappa(x,t)$ and $1/\rho(x,t)$ in (2.4) are both positive and bounded below away from zero, the time derivative $\kappa_{t}(x,t)$ of $\kappa(x,t)$ may change sign over time and thus either dampen or amplify the wave energy.
@@ 4 A time-dependent projection
denotes a time-dependent stiffness matrix weighted by $a(x,t)=1/\rho(x,t)$,
@@
the standard stifness-matrix, and
$\bm{B_{\kappa}}(t)=\big(\!\big(b(.,t)\psi_{k},\psi_{\ell}\big)\!\big)_{1\leq k,\ell\leq N}$
denotes a time-dependent mass matrix weighted by $b(x,t)=1/\kappa(x,t)$, whereas the right-hand side $\underline{F}(t)=\big(F_{1}(t),\ldots,F_{N}(t)\big)^{\top}$ is given by
@@
Since $\bm{B_{\kappa}}(t)$ and $\bm{K}$ are symmetric, while $\bm{K}$ is also positive definite, the $N\times N$ system of linear ordinary differential equations (4.3) has a unique solution $\underline{C}(t)$ if $\gamma K+B_{\kappa}(t)$ remains positive definite for all $t>0$. Hence for $\gamma>\Lambda\,(=\max\{0,-\lambda_{\min}({\bm{B}_{\kappa}}(t))/\lambda_{\min}({\bm{K}})\})$ sufficiently large, there exists a unique solution $\underline{C}(t)$ of (4.3) and therefore $w(t)$ is well-defined for $t\in J=(0,T)$; note that $\Lambda$ is bounded above independently of $h$ for shape-regular and quasi-uniform meshes.
For later analysis, we remark that (4.1) admits the following equivalent integral formulation

### v2

@@ 1 Introduction
Wave phenomena are central to a wide range of scientific and engineering applications in acoustics, electromagnetics or elasticity. Traditionally, numerical methods, be they finite element or finite difference methods, have primarily addressed wave propagation in passive (or static) media with material properties varying in space but not in time. Recent advances in material sciences, however, in particular the development of time-modulated metamaterials, have brought renewed attention to wave propagation in media whose properties vary in both space and time. Indeed, these engineered materials enable fine-grained, dynamic control over parameters such as refractive index and impedance, unlocking novel wave-matter interactions with no analogue in the static case [1]. From a mathematical perspective, these wave phenomena are modeled by second-order hyperbolic partial differential equations with coefficients that depend explicitly on both space and time, which introduces new analytical and computational challenges.
To this day, the Finite Element Method (FEM) remains probably the most versatile numerical method for the simulation of wave phenomena, as it naturally allows for high-order polynomial approximations, a key feature to efficiently represent the oscillatory nature of waves, but also accomodates complex geometry and discontinuous material interfaces. There is a long history of FEMs for second-order wave equations, starting in the 70’s with the seminal work of Dupont [9] and Baker [3, 2] who first proved optimal $H^{1}$- and $L^{2}$ convergence rates for standard conforming FEM. Later, convergence was proved in situations of low regularity [12], and more recently in polygons with corner singularities [15]. For non-conforming discontinuous Galerkin (DG) FEMs, convergence for the second-order wave equation was established during the past twenty years (e.g., [16, 11, 10, 5, 8]).
@@ 2 Wave equation in a time-varying medium
where $\mathcal{A}(t)$ is a self-adjoint, coercive and linear second-order elliptic partial differential and $\mathcal{B}(t)$ an arbitrary linear bounded operator. The well-posedness of the variational formulation of (2.5) follows from ([14], Theorem 1.1) under hypotheses about the temporal regularity of $\mathcal{A}(t)$ and $\mathcal{B}(t)$. Although $1/\kappa(x,t)$ and $1/\rho(x,t)$ in (2.4) are both positive and bounded below away from zero, the time derivative $\kappa_{t}(x,t)$ of $\kappa(x,t)$ will generally change sign over time and thus either dampen or amplify the wave energy.
@@ 4 A time-dependent projection
denotes a time-dependent stiffness matrix weighted by $a(x,t)=1/\rho(x,t)$
@@
the standard stifness-matrix, whereas
$\bm{B_{\kappa}}(t)=\big(b(.,t)\psi_{k},\psi_{\ell}\big)_{1\leq k,\ell\leq N}$
denotes a time-dependent mass matrix weighted by $b(x,t)=-\kappa_{t}(x,t)/\kappa^{2}(x,t)$, while the right-hand side $\underline{F}(t)=\big(F_{1}(t),\ldots,F_{N}(t)\big)^{\top}$ is given by
@@
Since $\bm{B_{\kappa}}(t)$ and $\bm{K}$ are symmetric, while $\bm{K}$ is also positive definite, the $N\times N$ system of linear ordinary differential equations (4.3) has a unique solution $\underline{C}(t)$ if $\gamma\bm{K+B_{\kappa}}(t)$ remains positive definite for all $t>0$. Hence for $\gamma>\Lambda\,(=\max\{0,-\lambda_{\min}({\bm{B}_{\kappa}}(t))/\lambda_{\min}({\bm{K}})\})$ sufficiently large, there exists a unique solution $\underline{C}(t)$ of (4.3) and therefore $w(t)$ is well-defined for $t\in J=(0,T)$; note that $\Lambda$ is bounded above independently of $h$ for shape-regular and quasi-uniform meshes.
For the subsequent analysis, we remark that (4.1) admits the following equivalent integral formulation: